40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 23 15:16:52 2017, gautam, Update, Optical Levers, ETMY Oplev HeNe needs to be replaced ETMY_Oplev.pdf
    Reply  Mon Jan 23 17:24:26 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is running hot ETMY_oplev__lasers.png
       Reply  Tue Jan 24 17:52:15 2017, Craig, Update, Optical Levers, ETMY Oplev HeNe is replaced 6x
          Reply  Wed Mar 1 08:28:04 2017, Steve, Update, Optical Levers, oplev laser RIN RIN_1103P.pngRIN_HeNe.png
             Reply  Thu Mar 2 13:59:04 2017, Steve, Update, Optical Levers, oplev laser RIN 1103P@2d.png
                Reply  Thu Mar 2 17:58:45 2017, rana, Update, Optical Levers, oplev laser RIN 
                   Reply  Fri Mar 3 17:24:21 2017, Steve, Update, Optical Levers, oplev laser RIN 3march17.pdf
                      Reply  Wed Mar 8 15:28:37 2017, Steve, Update, Optical Levers, oplev laser RIN RIN.jpgRIN_1103P_rotated.pdf
                      Reply  Mon Apr 10 09:49:32 2017, Steve, Update, Optical Levers, oplev laser RIN test planning 
                         Reply  Mon Apr 10 14:21:57 2017, rana, Update, Optical Levers, oplev laser RIN test planning 
       Reply  Tue May 30 13:36:58 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is replaced oplevETMY120d.png
          Reply  Tue Jan 15 18:16:00 2019, gautam, Update, Optical Levers, ETMY Oplev HeNe needs replacement OLRIN.pdfOLsums.png
             Reply  Fri Jan 18 12:52:07 2019, gautam, Update, Optical Levers, BS/PRM Oplev HeNe replaced OLRIN_20190118.pdf
Message ID: 12746     Entry time: Mon Jan 23 15:16:52 2017     Reply to this: 12747
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: ETMY Oplev HeNe needs to be replaced 

On the control room monitors, I noticed that the IR TEM00 spot was moving around rather a lot in the Y arm. The last time this happened had something to do with the ETMY Oplev, so I took a look at the 30 day trend of the QPD sum, and saw that it was decaying steeply (Steve will update with a long term trend plot shortly). I noticed the RIN also seemed rather high, judging by how much the EPICS channel reading for the QPD sum was jumping around. Attached are the RIN spectra, taken with the OL spot well centered on the QPD and the arms locked to IR. Steve will swap the laser out if it is indeed the cluprit.

Attachment 1: ETMY_Oplev.pdf  35 kB  | Hide | Hide all
ETMY_Oplev.pdf
ELOG V3.1.3-